OpenStudy (mikezack123):

For what value(s) of x is the quantity of 2x plus 3 all over the quantity of 5x plus 3 equals 2 over 5

OpenStudy (anonymous):

Is the question worded like that or are you just trying to make it tougher for us?

OpenStudy (mikezack123):

@iambatman you don't have to be rude wtheck...

OpenStudy (anonymous):

Not being rude but it's easier to write + and - signs then writing it in words.

OpenStudy (anonymous):

than*

OpenStudy (anonymous):

And is it multiply choice, because no solutions exist, if I'm reading your question right.

OpenStudy (mikezack123):

@iambatman its asking for the value of x

OpenStudy (mikezack123):

i think its no real numbers

OpenStudy (anonymous):

None, there's no solutions.

OpenStudy (anonymous):

\[5(2x+3)=2(5x+3) =>10x+15=10x+6=> 0=-9\]

OpenStudy (anonymous):

C

OpenStudy (mikezack123):

so it would be C.

OpenStudy (mikezack123):

yes thank you!

OpenStudy (anonymous):

Np